• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. McCormick

    Log entry time 20:16:24 on October 7,2002

    Entry number 87575

    keyword=Left Arm Dipole Quench

    The left arm dipole quenched around 6:30 p.m. with a Helium level and Lead A Voltage fault. It reset fine, but afterwards the magnet wouldn't ramp up. We contacted Ed and Mark and Javier for advice, and they suggested rebooting the IOC for the magnets. We did that, and after the reboot the current regulation on the left arm started working again. It took the left arm a long time to lock, and the NMR
    is very noisy, but now things seem to be working again.